i need help with math. a company lost 1/7 of its employees. if 132 of its employees were left, how many employees were there originally
Answer by stanbon(75887)

a company lost 1/7 of its employees. if 132 of its employees were left, how many employees were there originally
----------------------
Let original number of employees be "x":
Equation:
x - (1/7)x = 132
---
(6/7)x = 132
----
x = (7/6)132
----
x = 154
============
Cheers,
Stan H.
============